OpenStudy (anonymous):

Use the system to answer the question. 4x - 3y = 26 3x + 2y = 11 What is the x-coordinate of the solution to the system? (Points : 4) -3 1 5 7

OpenStudy (anonymous):

do you know how to solve this using substitution or elimination method?

OpenStudy (anonymous):

wouldnt you substitute the number in like 4(-3) - 3(0) ?

OpenStudy (anonymous):

well yeah you could do that for each choices and check whether that choice satisfy the equations. would you like to apply that instead of solving it?

OpenStudy (anonymous):

Yea. Can I do it then check back with you to see if my answer is correct?

OpenStudy (anonymous):

It doesnt seem like any of those choices work

OpenStudy (wolf1728):

Solving both equations: A) 4x - 3y = 26 B) 3x + 2y = 11 Multiply equation B by 1.5 B) 4.5x + 3y = 16.50 Then add this to A) A) 4x - 3y = 26 8.5x = 42.5 x = ??

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@wolf1728 could you help with another?

OpenStudy (wolf1728):

Well okay.

OpenStudy (anonymous):

A line has a slope of -6 that passes through the point (-12, 5). What is the equation of the line? (Points : 4) 6x + y = -77 6x + y = -67 -6x + y = 77 -6x + y = 17

OpenStudy (wolf1728):

b = y -m*x b = 5 -(-6*-12) b = 5 -72 b = -67 y = -6x -67 So looks like answer is 6x + y = -67

OpenStudy (anonymous):

thank you

OpenStudy (wolf1728):

u r welcome :-)